# A Representation Learning Approach to Feature Drift Detection in Wireless Networks

### v1

@@ Abstract
Artificial Intelligence (AI) is foreseen to be a centerpiece in next generation wireless networks enabling enabling ubiquitous communication as well as new services. However, in real deployment, feature distribution changes may degrade the performance of AI models and lead to undesired behaviors. To counter for undetected model degradation, we propose ALERT; a method that can detect feature distribution changes and trigger model re-training that works well on two wireless network use cases: wireless fingerprinting and link anomaly detection. ALERT includes three components: representation learning, statistical testing and utility assessment. We rely on Multi-layer Perceptron (MLP) for designing the representation learning component, on Kolmogorov–Smirnov (KS) and Population Stability Index (PSI) tests for designing the statistical testing and a new function for utility assessment. We show the superiority of the proposed method against ten standard drift detection methods available in the literature on two wireless network use cases.
Index Terms:
feature drift detection, machine learning, artificial intelligence, wireless networks, fingerprinting, link anomaly detection
@@ I Introduction
AI is foreseen to be a centerpiece in next generation wireless networks, including 6th Generation Wireless Cellular Networks (6G) and beyond cellular networks [1] by enabling enabling ubiquitous communication, new services including high-accuracy localization [2], anomaly fault and detection [3] as well as replacing traditionally networking functionality by AI based realization towards so-called AI native functionality [4, 5]. AI models are typically developed offline by using a pre-defined amount of data and a set of Machine Learning (ML) techniques that are tuned (semi-)manually to find the best performing combination for the respective training data [6]. However, when deployed in a real, production environment, the model development workflow is managed by the so-called AI/ML workflow [7] realized through Machine Learning Operations (MLOps) tools  [8]. The combination of those tools and their deployment enable MLOps pipelines that automatically manage the data preparation, model training, evaluation, selection and serving in production systems.
@@
To detect and signal distribution changes that may degrade the performance of AI in production, libraries able to detect data drifts while integrating with existing MLOps tools have been developed [12]. These libraries incorporate several drift detectors, defined as methods that observe a stream of data over time and determine for every new data point if the current distribution of the data has changed compared to a reference data set [13]. Several drift detection techniques as part of three such libraries have been recently benchmarked on two use-cases: occupancy detection and prediction of energy consumption [14]. To date, the only investigations of the drift phenomenon on wireless data are available in [10] including a Kolmogorov-Smirnov based detection technique and [11] that considered Isolation Forests and threshold to detect drifts in an illustrative example.
@@
  • Validation on two wireless Use Cases that utilise real-world data. We show that the ALERT method outperforms all the baseline models, achieving an overall F1-score of $0.9$ in the fingerprinting use case and $0.88$ in the links use case.
@@ II Related work
Data and concept drift are sometimes interchangeably used in the literature while in some cases one is considered as a superset of the other. In this paper we follow the definition from [15] where data drifts are categorized into four primary types: covariate drift also referred to a feature drift in this paper, prior probability drift, concept drift, and dataset shift. We group related works in three categories: works that develop new or analyze existing drift techniques, works that develop drift detection tools and systems and works that focus on studying specific use-cases.
@@ II-A Drift Detection Techniques
The methods from the third group, that monitor distributions on two different time-windows, compare the two distributions computed over the two windows using statistical tests and signal a drift when the distributions are not equal. The Kullback-Leibler (KL) divergence test, the PSI, a variation of the KL [18], as well as ADaptive WINdowing (ADWIN) all fall under this category [16]. Some other statistical tests as follow probably also fall in this group. The Energy Distance (ED) [19] that computes a statistical distance between two probability distributions, the Earth Mover’s Distance (EMD) [20] that computes the minimal cost that must be paid to transform one distribution into the other, the KS the quantifies the distance between the empirical distribution function of the sample and the cumulative distribution function of the reference distribution, or between the empirical distribution functions of two samples, the Kuiper test that is closely related to KS, etc [21].
@@ II-B Drift Detection Tools and Systems
The authors in [14] proposed D3Bench, a benchmarking too that enables the functional and non-functional evaluation of drift detection tools. In their analysis they benchmark three tools three open source tools for drift detection and found that Evidently AI stands out for its general data drift detection, whereas NannyML excels at pinpointing the precise timing of shifts and evaluating their consequent effects on predictive accuracy. The authors of [22] start from the observation that observe that not all data drifts lead to degradation in prediction accuracy and propose a new strategy which, using decision trees, is able to precisely pinpoint low-accuracy zones within ML models. The work triggers model improvement through active learning only in cases of harmful drifts that detrimentally affect model performance. Rather than triggering model retraining when drift is noticed, [23] proposed Matchmaker, a tool that dynamically identifies the batch of training data that is most similar to each test sample, and uses the ML model trained on that data for inference.
@@ III Drift Definition
    The covariate shift [15] or drift [24], also referred to as source 1 concept drift in [17], data drift in [26], is observed when $P_{j}(X)\neq P_{j+1}(X)$, while $P_{j}(y|X)=P_{j+1}(y|X)$. In such cases, the feature distribution changes, when new data $\{d_{j+1}\}$ are entered into the $D_{i}$ dataset, thus the reason we refer to covariate drift also as feature drift in this paper.
@@ IV Problem Statement
In this paper we focus on feature drift as defined in Section III and assume a model $M_{0}$ is trained on a dataset $D_{0}$, which we call the original training dataset. Given a new dataset $D_{1}$, our goal is to: (i) asses the existence of the feature drift phenomenon; (ii) estimate its effects on the $M_{0}$ model performance and (iii) decide whether the $M_{0}$ model should be re-trained with the $D_{1}$ dataset in order to increase its quality.
@@
Figure 1: An example of data drift between winter and spring data in the fingerprinting dataset.
Figure 2: An example of data drift between Normal and Anomalous wireless links in the links dataset.
@@ V Feature drift detection using the ALERT method
Figure 3: The proposed ALERT methodology for assessing the data drift given 2 datasets $D_{0}$ and $D_{1}$.
@@ V-A Design of the Representation Learning Component
We employ a supervised approach to learn the representation of $D_{0}$ through a lightweight MLP rather then relying on more complex and computationally expensive and data-hungry Convolutional Neural Network (CNN) or transformer-based architectures.
@@ V-B Design of the Statistical Testing Component
For the statistical testing depicted in a blue box on the lower side of Figure 3, we opt to use 2 well-established methods (KS [29, 30] and PSI) [18] instead of relying on a single one. This decision aims to cancel out the weakness of KS and PSI by aggregating their outcomes. It also provides a more stable output of the system which can be easily interpreted by decision-making mechanisms. More specifically, KS requires a large number of samples to produce accurate results and sometimes struggles with non-normal distributions [31]. On the other hand, it performs well when investigating larger datasets and does not require any information related with the probability density functions (Probability Density Function (PDF) s) of the involved datasets [32]. As for PSI, it can perform well, even if a lower amount of data is investigated [33] and is considered a stable and robust measure to assess the difference between two data distributions. Under this premise, ALERT computes the following statistical tests:
@@ V-C Design of the Utility Assessment Component
$U\in(0,1)$ encapsulates the finality utility we obtain if the model retraining action is selected, given the datasets $D_{0}$ and $D_{1}$. The expected utility $U$ considers the outputs of KS and KL tests to evaluate the statistical difference of the data $D_{0}$ and $D_{1}$ and the extracted features $R_{0}$ and $R_{1}$. As a result it contains valuable information on the model retraining operation which can be leveraged for decision-making.
@@ VI-A Dataset Description
The fingerprinting dataset [27] contains received signal strength (RSS) measurements made with Bluetooth Low Energy (BLE) technology, measured in dBm. The dataset consists of $505.000$ data points, organised over $23$ classes that represent 2D coordinates, which are collected during the spring and during the winter. It can be used for outdoor fingerprint-based localization applications, as presented in [27]. We organise the data into $31$ smaller datasets each containing $16.290$ samples, as follows: The first dataset $D_{0}$ is used to train a random forest classifier, as our $M_{0}$ model. For this reason, we refer to the $D_{0}$ as the original dataset, as depicted in Figure 3. Then we edit the rest of the datasets $D_{1}$ $D{30}$ so that each third dataset contains drifted data (i.e., $D_{2}$, $D_{5}$, $D_{8}$, $D_{11}$, $D_{14}$ etc.) while the rest of them (i.e,$D_{3}$, $D_{4}$, $D_{6}$, $D_{7}$, $D_{9}$ etc.) contain data from the same distribution as $D_{0}$. This is done to simulate various scenarios of data drifts which are encountered in different time frames i.e., in different versions of datasets collected in the field.
@@
We proceed to testing the $M_{0}$ model’s Macro Precision, Macro Recall and Macro F1-score with each dataset. As illustrated in Figure 4, the performance of the $M_{0}$ drops significantly when it is tested with a dataset that contains drifted samples compared to the $D_{0}$. This experimental set-up shows the existence of the drift and enables it’s detection through ALERT and selected baselines.
@@
Figure 4: The performance of the $M_{0}$ model with different fingerprinting datasets. The performance drops when the data drift phenomenon is present, since the $M_{0}$ is trained with the $D_{0}$ dataset.
The links dataset contains data from $8492$ timeseries, each one of which having $302$ data samples. The dataset is labeled and data is organized into $5$ anomaly classes. We split the data into $9$ smaller datasets each one containing $943$ timeseries, as follows: the first dataset $D_{0}$ is used to train a random forest classifier, as our $M_{0}$ model. Similarly to our approach when using the fingerprinting case, the $D_{0}$ is again the original dataset, which is illustrated in Figure 3. Then we split the rest of the datasets $D_{1}$ $D_{8}$ so that the $D_{1}$, $D_{2}$ and $D_{3}$ to contain samples from the same distribution as $D_{0}$ and the $D_{4}$, $D_{5}$, $D_{6}$, $D_{7}$ and $D_{8}$ to contain drifted samples. Again, we test the $M_{0}$ model’s Macro Precision, Macro Recall and Macro F1-score with each dataset and we present our finding in Figure 5. We observe that the performance of the $M_{0}$ drops when tested with drifted samples, similarly to the fingerprinting model.
@@
Figure 5: The performance of the $M_{0}$ model with different links datasets. The model performance deteriorates when the $M_{0}$ is tested with datasets that contain data drifts.
@@ VI-C Baseline Selection
We compare the ALERT technique with state-of-the art methods that exist in the literature. We choose two broader types of methods for comparison, namely statistical methods and distance-based methods. Statistical methods leverage statistical indexes (such as mean values, sampling variations and variance) to predict data drifts. In this work, we formulate a baseline using the following statistical methods: (i) Kuiper test [34]; (ii) Cramer-Von Mises (CVM) [35]; (iii) Welch Test (WelchT) [36]; (iv) Chi Square test [37]; (v) Mann-Whitney U test (Mann Whitney) [38]; (vi) Andreson Darling Test [39]; and (vii) Kolmogorov-Smirnov (KS) test [40]. On the contrary, distance-based methods focus on estimating the distance between two data distributions by measuring the dissimilarity between them through distance functions. In this work we use as a baseline the following distance-based methods: (i) Population Stability Index (PSI) [41]; (ii) Energy Distance [42]; and (iii) Earth Mover’s Distance (EMD) [43].
@@ VI-D Training and Evaluation
We use the $D_{1}$ $D_{30}$ stemming from the fingerprinting and the $D_{1}$ $D_{8}$ stemming from the links datasets to evaluate ALERT and to assess whether it is able to identify the data drift phenomenon accurately. For the implementation of ALERT, we use the python programming language; whereas the code base for the Baseline methods is provided by [44]. For each validation Use Case, we train the ALERT method’s MLP for $3$ epochs using the $D_{0}$ dataset and then, we utilize the trained model to assess the existence of the data drift phenomenon.

### v2

@@ Abstract
Artificial Intelligence (AI) is foreseen to be a centerpiece in next generation wireless networks enabling ubiquitous communication as well as new services. However, in real deployment, feature distribution changes may degrade the performance of AI models and lead to undesired behaviors. To counter for undetected model degradation, we propose ALERT; a method that can detect feature distribution changes and trigger model re-training that works well on two wireless network use cases: wireless fingerprinting and link anomaly detection. ALERT includes three components: representation learning, statistical testing and utility assessment. We rely on Multi-layer Perceptron (MLP) for designing the representation learning component, on Kolmogorov–Smirnov (KS) and Population Stability Index (PSI) tests for designing the statistical testing and a new function for utility assessment. We show the superiority of the proposed method against ten standard drift detection methods available in the literature on two wireless network use cases.
@@ I Introduction
Artificial Intelligence (AI) is foreseen to be a centerpiece in next generation wireless networks, including 6th Generation Wireless Cellular Networks (6G) and beyond cellular networks [1] by enabling ubiquitous communication, new services including high-accuracy localization [2], anomaly fault and detection [3] as well as replacing traditionally networking functionality by AI based realization towards so-called AI native functionality [4, 5]. AI models are typically developed offline by using a pre-defined amount of data and a set of Machine Learning (ML) techniques that are tuned (semi-)manually to find the best performing combination for the respective training data [6]. However, when deployed in a real, production environment, the model development workflow is managed by the so-called AI/ML workflow [7] realized through Machine Learning Operations (MLOps) tools  [8]. The combination of those tools and their deployment enable MLOps pipelines that automatically manage the data preparation, model training, evaluation, selection and serving in production systems.
@@
To detect and signal distribution changes that may degrade the performance of AI in production, libraries able to detect feature drifts while integrating with existing MLOps tools have been developed [12]. These libraries incorporate several drift detectors, defined as methods that observe a stream of data over time and determine for every new data point if the current distribution of the data has changed compared to a reference data set [13]. Several drift detection techniques as part of three such libraries have been recently benchmarked on two use-cases: occupancy detection and prediction of energy consumption [14]. To date, the only investigations of the drift phenomenon on wireless data are available in [10] including a Kolmogorov-Smirnov based detection technique and [11] that considered Isolation Forests and threshold to detect drifts in an illustrative example.
@@
  • Validation on two wireless Use Cases that utilize real-world data. We show that the ALERT method outperforms all the baseline models, achieving an overall F1-score of $0.9$ in the fingerprinting use case and $0.88$ in the links use case.
@@ II Related work
Data and concept drift are sometimes interchangeably used in the literature while in some cases one is considered as a superset of the other. In this paper we follow the definition from [15] where feature drifts are categorized into four primary types: covariate drift also referred to a feature drift in this paper, prior probability drift, concept drift, and dataset shift. We group related works in three categories: works that develop new or analyze existing drift techniques, works that develop drift detection tools and systems and works that focus on studying specific use-cases.
@@ II-A Drift Detection Techniques
The methods from the third group, that monitor distributions on two different time-windows, compare the two distributions computed over the two windows using statistical tests and signal a drift when the distributions are not equal. The Kullback-Leibler (KL) divergence test, the Population Stability Index (PSI), a variation of the KL [18], as well as ADaptive WINdowing (ADWIN) all fall under this category [16]. Some other statistical tests as follow probably also fall in this group. The Energy Distance (ED) [19] that computes a statistical distance between two probability distributions, the Earth Mover’s Distance (EMD) [20] that computes the minimal cost that must be paid to transform one distribution into the other, the Kolmogorov-Smirnov test (KS) the quantifies the distance between the empirical distribution function of the sample and the cumulative distribution function of the reference distribution, or between the empirical distribution functions of two samples, the Kuiper test that is closely related to KS, etc [21].
@@ II-B Drift Detection Tools and Systems
The authors in [14] proposed D3Bench, a benchmarking tool that enables the functional and non-functional evaluation of drift detection tools. In their analysis they benchmark three open source tools for drift detection and found that Evidently AI stands out for its general feature drift detection, whereas NannyML excels at pinpointing the precise timing of shifts and evaluating their consequent effects on predictive accuracy. The authors of [22] start from the observation that observe that not all feature drifts lead to degradation in prediction accuracy and propose a new strategy which, using decision trees, is able to precisely pinpoint low-accuracy zones within ML models. The work triggers model improvement through active learning only in cases of harmful drifts that detrimentally affect model performance. Rather than triggering model retraining when drift is noticed, [23] proposed Matchmaker, a tool that dynamically identifies the batch of training data that is most similar to each test sample, and uses the ML model trained on that data for inference.
@@ III Drift Definition
    The covariate shift [15] or drift [24], also referred to as source 1 concept drift in [17], feature drift in [26], is observed when $P_{j}(X)\neq P_{j+1}(X)$, while $P_{j}(y|X)=P_{j+1}(y|X)$. In such cases, the feature distribution changes, when new data $\{d_{j+1}\}$ are entered into the $D_{i}$ dataset, thus the reason we refer to covariate drift also as feature drift in this paper.
@@ IV Problem Statement
In this paper, we focus on feature drift as defined in Section III and assume a model $M_{0}$ is trained on a dataset $D_{0}$, which we call the original training dataset. Given a new dataset $D_{1}$, our goal is to: (i) asses the existence of the feature drift phenomenon; (ii) estimate its effects on the $M_{0}$ model performance and (iii) decide whether the $M_{0}$ model should be re-trained with the $D_{1}$ dataset in order to increase its quality.
@@
Figure 1: An example of feature drift between winter and spring data in the fingerprinting dataset.
Figure 2: An example of feature drift between Normal and Anomalous wireless links in the links dataset.
@@ V Feature drift detection using the ALERT method
Figure 3: The proposed ALERT methodology for assessing the feature drift given 2 datasets $D_{0}$ and $D_{1}$.
@@ V-A Design of the Representation Learning Component
We employ a supervised approach to learn the representation of $D_{0}$ through a lightweight Multi-Layer Perceptron (MLP) rather than relying on more complex and computationally expensive and data-hungry Convolutional Neural Network (CNN) or transformer-based architectures.
@@ V-B Design of the Statistical Testing Component
For the statistical testing depicted in the blue box on the lower side of Figure 3, we opt to use $2$ well-established methods (KS [29, 30] and PSI) [18], instead of relying on a single one. This decision aims to: (i) reduce the input parameters of ALERT, as much as possible; (ii) cancel out the weakness of KS and PSI by aggregating their outcomes; (iii) reinforce ALERT with the capacity of detecting both small and medium distribution drifts; (iv) introduce sampling symmetry to ALERT; and (v) stabilize and simplify ALERT’s outputs so that to be easily usable and interpretable by decision-making mechanisms.
Rather than relying on pre-set input parameters or prior knowledge regarding the (Probability Density Function (PDF) s), KS is able to internally estimate the distribution characteristics of the input data. As a result, KS swiftly detects small distribution changes (even in large datasets) and produces a p-value that can be easily interpreted by external systems [31]. However, KS is prone to false positives and is asymmetric thus, if the user swaps the baseline and sample distributions it will produce different outcomes. Also, it requires a large number of samples to accurately function and may struggle with non-normal distributions [32]. PSI, on the other hand, can perform well, even if a lower amount of data is used [33] and is considered a stable and robust measure to assess drifts even when non-normal distributions are considered. Also, PSI produces interpretable outputs, is very good at identifying moderate or larger drifts, and is fully symmetric. On the flip side, it requires the data binning parameter to be set externally, which separates the input data into predefined intervals. The combination of KS and PSI provides a unified framework that cancels out the weaknesses of each method and increases the robustness of ALERT. The importance of both KS and PSI is further discussed in Section VII-E, in which we perform an ablation study and we showcase the individual contributions of KS and PSI to the overall ALERT’s utility score. There, it is evident that if the KS or the PSI component is removed, the performance of ALERT drops by a large margin. To accomplish the unification of KS and PSI, ALERT computes the following statistical tests:
@@ V-C Design of the Utility Assessment Component
$U\in(0,1)$ encapsulates the final utility we obtain if the model retraining action is selected, given the datasets $D_{0}$ and $D_{1}$. The expected utility $U$ considers the outputs of KS and KL tests to evaluate the statistical difference of the data $D_{0}$ and $D_{1}$ and the extracted features $R_{0}$ and $R_{1}$.
@@ VI-A Dataset Description
TABLE I: The details of the drifted and non-drifted datasets which are created based on the fingerprinting and links.
fingerprinting links
Data types BLE measurements Time series
Total samples $505.000$ $8.493\times 302$
# Classes $25$ 55
$M_{0}$ training $D_{0}$ $D_{0}$
$D_{3},D_{6},D_{9},D_{12},$ $D_{4},D_{5},D_{6},$
Drifted data $D_{15},D_{18},D_{21},D_{24},$ $D_{7},D_{8}$
$D_{27},D_{30}$
$D_{1},D_{2},D_{4},D_{5},$
$D_{7},D_{8},D_{10},D_{11},$
Non-drifted data $D_{13},D_{14},D_{16},D_{17},$ $D_{1},D_{2},D_{3}$
$D_{19},D_{20},D_{22},D_{23},$
$D_{25},D_{26},D_{28},D_{29}$
Cause of drift Spring data is New anomalies are
mixed with winter data added to the data
@@
The fingerprinting dataset [27] contains received signal strength (RSS) measurements made with Bluetooth Low Energy (BLE) technology, measured in dBm. The dataset consists of $505.000$ data points, organised over $25$ classes that represent 2D coordinates, which are collected during the spring and during the winter, as described in Section IV. It can be used for outdoor fingerprint-based localization applications, similarly to [27]. We organise the data into $31$ smaller datasets each containing $16.290$ samples, as follows: The first dataset $D_{0}$, which contains samples collected during the winter season only, is used to train a random forest classifier, as our $M_{0}$ model. For this reason, we refer to the $D_{0}$ as the original dataset, as depicted in Figure 3. Then we edit the rest of the datasets $D_{1}$ - D30D{30} so that each third dataset contains samples stemming from spring measurements, which are drifted (i.e., $D_{3}$, $D_{4}$, $D_{6}$, $D_{7}$, $D_{9}$ etc.). The rest of the datasets (i.e., $D_{2}$, $D_{5}$, $D_{8}$, $D_{11}$, $D_{14}$ etc.) contain samples that are collected during winter and belong to the same distribution with $D_{0}$. Through this process, we can simulate various scenarios of feature drifts which are encountered in different time frames i.e., in different versions of datasets collected in the field. Table I summarizes the drift creation details for the fingerprinting dataset.
@@
In the sequel, we test the $M_{0}$ model’s Macro Precision, Macro Recall and Macro F1-score with each created dataset ($D_{1}-D_{30}$). As illustrated in Figure 4, the performance of the $M_{0}$ drops significantly when it is tested with a dataset that contains drifted samples. This experimental set-up shows the existence of the drift and enables its detection through ALERT and selected baselines.
@@
Figure 4: The performance of the $M_{0}$ model with different fingerprinting datasets. The performance drops when the feature drift phenomenon is present, since the $M_{0}$ is trained with the $D_{0}$ dataset.
@@
Figure 5: The performance of the $M_{0}$ model with different links datasets. The model performance deteriorates when the $M_{0}$ is tested with datasets that contain feature drifts.
The links dataset contains data from $8492$ timeseries, each one of which having $302$ data samples. The dataset is labeled and is organized into $5$ anomaly classes. We split the data into $9$ smaller datasets each one containing $943$ timeseries, as follows: the first dataset $D_{0}$ is used to train a random forest classifier, as our $M_{0}$ model. Similarly to our approach when using the fingerprinting case, the $D_{0}$ is again the original dataset, which is illustrated in Figure 3. Then we split the rest of the datasets $D_{1}$ - $D_{8}$ so that the $D_{1}$, $D_{2}$ and $D_{3}$ to contain samples from the same distribution as $D_{0}$, while the $D_{4}$, $D_{5}$, $D_{6}$, $D_{7}$ and $D_{8}$ to contain drifted samples, from new anomaly types. Table I summarizes the details of this process. We then test the $M_{0}$ model’s Macro Precision, Macro Recall and Macro F1-score with each created dataset and we present our findings in Figure 5. We observe that the performance of the $M_{0}$ drops when used with drifted samples, similarly to the fingerprinting model.
@@ VI-C Baseline Selection
We compare the ALERT technique with state-of-the art methods that exist in the literature. We choose three broader types of methods for comparison, namely statistics-based methods, distance-based methods and ML-based methods. Statistical methods leverage statistical indexes (such as mean values, sampling variations and variance) to predict feature drifts. In this work, we formulate a baseline using the following statistical methods: (i) Kuiper test [34]; (ii) Cramer-Von Mises (CVM) [35]; (iii) Welch Test (WelchT) [36]; (iv) Chi Square test [37]; (v) Mann-Whitney U test (Mann Whitney) [38]; (vi) Andreson Darling Test [39]; and (vii) Kolmogorov-Smirnov (KS) test [40]. On the contrary, distance-based methods focus on estimating the distance between two data distributions by measuring the dissimilarity between them through distance functions. In this work we use as a baseline the following distance-based methods: (i) Population Stability Index (PSI) [41]; (ii) Energy Distance [42]; and (iii) Earth Mover’s Distance (EMD) [43]. ML-based methods utilize Deep Learning or Machine Learning models to either form an intermediate representation of the data under investigation, or to estimate the features of data distributions. In this paper we use the following ML-based methods as baselines: (i) Adversarial Detection (ML-AD) [44]; (ii) Embedding-based domain classification (ML-EDC) [45]; and (iii) Data-driven deep Density Estimation (ML-DDE) [46].
@@ VI-D Training and Evaluation
We use the $D_{1}$ - $D_{30}$ stemming from the fingerprinting and the $D_{1}$ - $D_{8}$ stemming from the links datasets to evaluate ALERT and to assess whether it is able to identify the feature drift phenomenon accurately. For the implementation of ALERT, we use the python programming language; whereas the code base for the Baseline methods is provided by [44], [45], [46] and [47]. For each validation Use Case, we train the ALERT method’s MLP for $3$ epochs using the $D_{0}$ dataset and then, we utilize the trained model to assess the existence of the feature drift phenomenon.
Since the distance-based and statistics-based methods are provided in a ready-to-use form in GitHub repositories [47], their evaluation is trivial. On the other hand, ML-based methods require some adaptations in order to properly work with our datasets. For this reason, we have performed the following adaptations:
For the ML-AD method
we use the open source code provided by the author [44] and we opt to train a ResNet-28-10 model from scratch to fit the $D_{0}$ dataset. Then, we fine-tune the model using the provided Bayesian ensemble method, which utilizes one optimizer (SGD) and one regularizer (Prior Regularizor) in parallel. In the sequel, we save the trained model and test its performance with each new dataset ($D_{1}$ - $D_{30}$ for the fingerprinting data, or $D_{1}$ - $D_{8}$ for the links data). This approach essentially performs adversarial detection for each new dataset. The outcomes of each test determine if the drift phenomenon is detected. Thus, when the model detects adversarial samples, we assess that the drift phenomenon is present.
For the ML-EDC method
we use an unsupervised representation learning approach for drift detection. More specifically, we train the DriftLens model [45] with our $D_{0}$ dataset in order for the model to learn the distribution properties of the input data. Then, we utilise the $D_{1}$ dataset to estimate the threshold distance values that discriminate between drift and no-drift conditions, as suggested by the authors. Afterwards, we utilise the rest of the datasets to perform inference of the trained model.
For the ML-DDE method
we first train a new model to recognize the Probability Density Function (PDF) characteristics of our $D_{0}$ dataset. For the fingerprinting dataset we use a $2$-Dimensional data structure ($X$ and $Y$ coordinates), while for the links dataset we use a $5$-Dimensional data structure ($1$ dimension for each class). In the sequel we conduct inference using the trained model and the rest of the datasets. In order to assess the presence of drift phenomena, we use the KL-divergence metric, as suggested by the authors [46], to measure the distance between the estimated PDF of the $D_{0}$ and the estimated PDF of the dataset under investigation.